OpenStudy (fantalux):

Which is the MOST simplified version of the ratio 27:81 ?

OpenStudy (anonymous):

27/81 = 1/3 27:81 = 1:3

OpenStudy (fantalux):

ty!

OpenStudy (lizz123):

1/3